40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 10:52:38 2019, gautam, Update, SUS, EY door opened 
    Reply  Tue Jan 15 15:49:47 2019, gautam, Update, SUS, EY door opened EY_OSEMs.png
       Reply  Sat Jan 19 15:33:18 2019, gautam, Update, SUS, ETMY OSEMs faulty Screen_Shot_2019-01-19_at_3.32.35_PM.pngETMY_sensors_1231832635.pdf
          Reply  Tue Jan 22 20:36:53 2019, gautam, Update, SUS, ETMY OSEMs faulty 
             Reply  Wed Jan 23 12:39:18 2019, gautam, Update, SUS, EY chamber work 6x
                Reply  Wed Jan 23 23:12:44 2019, gautam, Update, SUS, Prep for FC cleaning IY_level_before.pdf
                   Reply  Thu Jan 24 15:32:31 2019, gautam, Update, SUS, Y arm cavity side first contact applied Yarm_FC.pdfOSEMfilter.png
                      Reply  Tue Jan 29 22:12:40 2019, gautam, Update, SUS, Alignment prep 
                         Reply  Wed Jan 30 11:54:24 2019, gautam, Update, SUS, More alignment prep IMG_5958.JPGIMG_5962.JPG
                            Reply  Wed Jan 30 19:25:40 2019, gautam, Update, SUS, Xarm cavity alignment Xarm.pngXarm_IR.png
                               Reply  Fri Feb 1 01:24:06 2019, gautam, Update, SUS, Almost ready for pumpdown tomorrow 
                                  Reply  Fri Feb 1 13:16:50 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                     Reply  Fri Feb 1 21:52:57 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                  Reply  Fri Feb 1 14:44:14 2019, gautam, Update, SUS, Y arm FC cleaning and reinstall IMG_5974.JPGIMG_5986.JPGIMG_5992.JPG
    Reply  Wed Jan 16 16:25:25 2019, gautam, Update, SUS, Yarm locked Yarm_locked.png
       Reply  Fri Jan 18 21:34:18 2019, gautam, Update, SUS, Unused optic on EY table IMG_5408.JPG
          Reply  Sat Jan 19 05:07:45 2019, Koji, Update, SUS, Unused optic on EY table 
Message ID: 14416     Entry time: Thu Jan 24 15:32:31 2019     In reply to: 14415     Reply to this: 14422
Author: gautam 
Type: Update 
Category: SUS 
Subject: Y arm cavity side first contact applied 

EY:

  • A clean cart was setup adjacent to the HEPA-enclosed mini cleanroom area (it cannot be inside the mini cleanroom, because of lack of space). 
  • The FC tools (first contact, acetone, beakers, brushes, PEEK mesh, clean scissors, clean tweezers, Canon camera, green flashlight) were laid out on this cart for easy access.
  • I inspected the optic - the barrel had a few specks of dust, and the outer 1.5" annular region of the HR face looked to have some streak marks
    • I was advised not to pre-wipe the HR side with any solvents
    • The FC was only applied to the centran ~1-1.5" of the optic
  • After applying the FC, I spent a few minutes inspecting the status of the OSEMs 
    • Three out of the four face OSEMs, as well as the side OSEM, did not have a filter in
    • I inserted filters into them.
  • Closed up the chamber with light door, left HEPA unit on and the mini cleanroom setup intact for now. We will dismantle everything after the pumpdown.

IY:

  • Similar setup to EY was implemented
  • Removed side OSEM from ITMY.
  • Double-checked that EQ stops were engaged.
  • Moved the OSEM cable tower to free up some space for accommodating ITMY.
  • Undid the clamps of ITMY, moved it to the NE corner of the IY table.
  • Inspected the optic - it was much cleaner than the 2016 inspection, although the barrel did have some specks of dust.
  • Once again, I applied first contact to the central ~1.5" of the HR surface.
  • Checked status of filters on OSEMs - this time, only the UL coil required a filter.
    • Attachment #3 shows the sensor voltage DC level before and after the insertion of the filter. There is ~0.1% change.
    • The filters were found in a box that suggests they were made in 2002 - but Steve tells me that it is just stored in a box with that label, and that since there are >100 filters inside that box, he thinks they are the new ones we procured in 2016. The coating specs and type of glass used are different between the two versions.

The attached photo shows the two optics with FC applied.

My original plan was to attempt to close up tomorrow. However, we are still struggling with Satellite box issues. So rather than rush it, we will attempt to recover the Y arm cavity alignment on Monday, satellite box permitting. The main motivation is to reduce the deadtime between peeling off the F.C and starting the pumpdown. We will start working on recovering the cavity alignment once the Sat box issues are solved.

Attachment 1: Yarm_FC.pdf  12.236 MB  | Hide | Hide all
Yarm_FC.pdf
Attachment 2: OSEMfilter.png  46 kB  Uploaded Sat Jan 26 13:56:30 2019  | Hide | Hide all
OSEMfilter.png
ELOG V3.1.3-